Suppose A and B are mutually exclusive events defined on the same sample space such that P(A)=0.37 and P(B)=c.


1. What is the maximum value of c?

2. What is P(A|B)?


1
Expert's answer
2021-10-04T16:24:19-0400

Solution:

(1):

For mutually exclusive, P(A)+P(B)=P(AUB) and P(A"\cap"B)=0

P(A)+P(B)=P(AUB)

"\Rightarrow" 0.37+c=1 [Since P(AUB)"\le" 1 and we need maximum value of c]

"\Rightarrow"c=0.63

(2):

"P(A|B)=\dfrac{P(A\cap B)}{P(B)}=\dfrac{0}{0.63}=0"